OpenStudy (anonymous):

The radius of the Milky Way Galaxy is about 50,000 light-years. Write this number in scientific notation. What is the radius of the Milky Way Galaxy in miles? Write your final answer in scientific notation and show at least one step of your work.

OpenStudy (anonymous):

Tully00, we are here to help you. However, we cannot make you homework. You have to try first, then if you have a question, we can help you. You have to try it first!

OpenStudy (anonymous):

I have tried to work them out. what i got for the first one was 5 X 10 4th power

OpenStudy (anonymous):

That is the correct answer. Try 0.000245 now

OpenStudy (anonymous):

2.45 X 10 -4th power

OpenStudy (anonymous):

exactly. That is the answer. I thing you are already done with scientist notation :) Do you have any question?

OpenStudy (anonymous):

yes the question that I was having trouble with was the second one. How do I find the radius

OpenStudy (anonymous):

in miles

OpenStudy (anonymous):

I need 2 minutes to get the answer. Wait please

OpenStudy (anonymous):

If the radius of the Milky Way Galaxy is about 50,000 (5x10^4) light-years, and 1 light-year is 5,878,630,000 (5.88x10^12), then: \[r=(5x10^{4} Light-years)*(5.88x10^{12} miles/1 light-year)\] then r=29.4x10^16 r=2.94x10^17 This is the answer

OpenStudy (anonymous):

Light-year in the numerator cancel with the numerator in the denominator, so you get the radius in miles

OpenStudy (anonymous):

got it ttully00?

OpenStudy (anonymous):

yeah it seems kind of complicated.

OpenStudy (anonymous):

what part you dont understand?

OpenStudy (anonymous):

how did you get the 29.4

OpenStudy (anonymous):

Multiplying 5x10^4 with 5.88x10^12. You have to multiply 5 and 5.88 first so you get 29.4. After that you have to multiply 10^4 and 10^12 = 10^4 (for this multiplication the base (10) remains and the exponents are added (4+12=16))

OpenStudy (anonymous):

There is a mistake in the last post. Read this instead :) Multiplying 5x10^4 with 5.88x10^12. You have to multiply 5 and 5.88 first so you get 29.4. After that you have to multiply 10^4 and 10^12 = 10^16 (for this multiplication the base (10) remains and the exponents are added (4+12=16))

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

O ok

OpenStudy (anonymous):

Try multiplying 5x10^5 and 2x10^8 :)